In a wire, when elongation is 2 cm energy stored is E. if it is stretched by 10 cm, then the energy stored will be

(a) E (b) 2 E
(c) 4 E (d) 25 E

Grade:Upto college level

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

The answer is option d

Regards
Vikas TU
3 years ago

We know, potential energy of spring is given by (1/2)kx² where k is spring constant and x is stretched distance.

Here , x=2cm so, U=(1/2)k*2²=2k

For, x=10cm :-

E=(1/2)k*10²=50k=25(2k)=25U.

Final potential energy will be 25U.